# Ill-posedness of the Euler equations and inviscid limit of the Navie-Stokes equations in Besov spaces

### v1

@@ 2 Preliminaries
$(\mathcal{F}u)(k)=\widehat{u}(k)=\int_{\mathbb{T}^{d}}e^{-\mathrm{i}x\cdot k}u(x)\mathrm{d}x.$
and the inverse Fourier transform $\mathcal{F}^{-1}_{\mathbb{T}^{d}}:\mathcal{S}(\mathbb{Z}^{d})\rightarrow\mathcal{D}(\mathbb{T}^{d})$ as
$(\mathcal{F}^{-1}u)(x)=\frac{1}{(2\pi)^{d}}\sum_{k\in\mathbb{Z}^{d}}u(k)e^{\mathrm{i}x\cdot k}.$
@@
$u(x)=\frac{1}{(2\pi)^{d}}\sum_{k\in\mathbb{Z}^{d}}\widehat{u}(k)e^{\mathrm{i}x\cdot k}.$
We are interested in solutions which take values in the Besov space $B_{p,r}^{s}(\mathbb{T}^{d})$. Recall that one way to define this space requires a dyadic partition of unity. Given a smooth bump function $\chi$ supported on the ball of radius $4/3$, and equal to 1 on the ball of radius $3/4$, we set $\varphi(\xi)=\chi(2^{-1}\xi)-\chi(\xi)$ and $\varphi_{j}(\xi)=\varphi(2^{-j}\xi)$. then we deduce that $\varphi$ satisfies that ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{T}^{d}:3/4\leq|\xi|\leq 8/3\right\}$ and $\varphi(\xi)\equiv 1$ for $24/3\leq|\xi|\leq 3/$ 2. Using this partition, we define the periodic dyadic blocks as follows
@@
$\Delta_{-1}u=\frac{1}{(2\pi)^{d}}\sum_{\xi\in\mathbb{Z}^{d}}\chi(\xi)\widehat{u}(\xi)e^{\mathrm{i}x\cdot\xi},$
$\Delta_{j}u=\frac{1}{(2\pi)^{d}}\sum_{\xi\in\mathbb{Z}^{d}}\varphi_{j}(\xi)\widehat{u}(\xi)e^{\mathrm{i}x\cdot\xi},\quad\text{ if }\quad j\geq 0.$
@@
Lemma 2.1
Let $3\leq m\in\mathbb{Z}$ and $-1\leq j\in\mathbb{Z}$, we have
$\Delta_{j}\left[\cos\left(\frac{11}{8}2^{m}x_{1}\right)\right]=\begin{cases}0,&\text{if}\quad j\neq m,\\ \cos(\frac{11}{8}2^{m}x_{1}),&\text{if}\quad j=m.\end{cases}$
Proof. We set $\lambda=\frac{11}{8}2^{m}$ and $\vec{\lambda}=(\lambda,0,\cdots,0)$ for simplicity. Notice that $\varphi_{j}(k)$ is symmetric, i.e., $\varphi_{j}(k)=\varphi_{j}(|k|)$, we deduce
$\Delta_{j}\left(\cos(\lambda x_{1})\right)$ $=\sum_{k\in\mathbb{Z}^{d}}\varphi_{j}(k)\left(\mathbf{1}_{\vec{\lambda}}(k)+\mathbf{1}_{-\vec{\lambda}}(k)\right)e^{\mathrm{i}x\cdot k}$
$=\sum_{\ell\in\mathbb{Z}}\varphi_{j}(\ell)\left(\mathbf{1}_{\lambda}(\ell)+\mathbf{1}_{-\lambda}(\ell)\right)e^{\mathrm{i}x_{1}\ell}$
@@
Thus we obtain the desired result of Lemma 2.1.
Remark 2.1
For large fixed $m$, $\varphi_{j}\left(\frac{11}{8}2^{m}x_{1}\right)$ vanish except for $j=m$ due to the support condition of $\varphi$ and $\varphi(k)\equiv 1$ for $24/3\leq|k|\leq 3/$ 2.
@@
where $f(x)$ is a bounded real-valued periodic function of one variable of the following form
@@
$u(t,x)=\left(1,f\left(x_{1}-t\right),0,\ldots,0\right),$ (3.3)
@@
where $g(x)$ is a bounded real-valued period function of one variable of the following form
@@
which is a period function which also satisfies the incompressible Euler equations (E) with initial condition $u_{0}(x)=(1,g(x_{1}),0,\cdots,0)$.
@@
5 Proof of Theorem 1.3
Let $\lambda_{n}=1+\frac{1}{n}$ and $f(x)$ be given by (3.2), we consider
$u(t,x)=\left(1,f(x_{1}-t),0,\ldots,0\right),$
$u_{n}(t,x)=\left(\lambda_{n},f\left(x_{1}-\lambda_{n}t\right),0,\ldots,0\right).$
It is not difficult to verify that both $u(t,x)$ and $u_{n}(t,x)$ are periodic solutions to the Euler equations (E) with initial data $u_{0}=(1,f(x_{1}),0,\cdots,0)$ and $u_{0,n}=(\lambda_{n},f(x_{1}),0,\cdots,0)$, respectively.
Difference of initial data.  Obviously, we see that
$\|u_{0,n}-u_{0}\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}=\frac{1}{n}\|1\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}\leq\frac{1}{n}\to 0,\quad\text{as}\;n\to\infty,$
where we have used $\Delta_{j}1=0$ if $j\neq-1$.
Difference of solutions.  Obviously,
$u_{n}(t,x)-u(t,x)=\left(\frac{1}{n},f(x_{1}-\lambda_{n}t)-f(x_{1}-t),0,\ldots,0\right).$
Next we shall prove that
$\|u_{n}(t,x)-u(t,x)\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}\nrightarrow 0,\quad\text{as}\;n\to\infty.$
A straightforward computation shows that
$f(x_{1}-\lambda_{n}t)-f(x_{1}-t)$ $=\sum\limits^{\infty}_{j=3}2^{-js}\left[\cos\left(\frac{11}{8}2^{j}x_{1}\right)\mathbf{c}_{j}(t)+\sin\left(\frac{11}{8}2^{j}x_{1}\right)\mathbf{d}_{j}(t)\right],$
where
$\mathbf{c}_{j}(t)=\cos\left(\frac{11}{8}2^{j}\lambda_{n}t\right)-\cos\left(\frac{11}{8}2^{j}t\right),$
$\mathbf{d}_{j}(t)=\sin\left(\frac{11}{8}2^{j}\lambda_{n}t\right)-\sin\left(\frac{11}{8}2^{j}t\right).$
Then, by Lemma 2.1, we have for some $n$ large enough
$2^{ns}\Delta_{n}[f(x_{1}-\lambda_{n}t)-f(x_{1}-t)]$ $=\cos\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{c}_{n}(t)+\sin\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{d}_{n}(t).$
Letting $\frac{11}{8}2^{n}t_{n}=n\pi$, then one has $\mathbf{c}_{n}(t_{n})=\pm 2$ and $\mathbf{d}_{n}(t_{n})=0$, which give that
$\quad 2^{ns}\left\|\Delta_{n}[f(x_{1}-\lambda_{n}t_{n})-f(x_{1}-t_{n})]\right\|_{L^{p}(\mathbb{T}^{d})}=2(2\pi)^{d-1}\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}(\mathbb{T})}.$
Thus, we have for some $n$ large enough
$\|u_{n}(t_{n},x)-u(t_{n},x)\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}$ $\geq\|f(x_{1}-\lambda_{n}t_{n})-f(x_{1}-t_{n})\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}-\frac{1}{n}$
$\geq 2^{ns}\left\|\Delta_{n}[f(x_{1}-\lambda_{n}t)-f(x_{1}-t_{n})]\right\|_{L^{p}(\mathbb{T}^{d})}-\frac{1}{n}$
$\geq 2c_{0}(2\pi)^{d-1}-\frac{1}{n}.$
Letting $n\to\infty$, we complete the proof of Theorem 1.3.
@@
$\Delta_{n}\left[f_{n}(t,x_{1})-f(x_{1}-t_{n})\right]$
$=$ $~{}2^{-ns}\left[\cos\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{e}_{n}(t)+\sin\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{f}_{n}(t)\right]\mathbf{g}_{n}(t),$
@@
[…]
$\liminf_{\varepsilon_{n}\to 0^{+}}\left\|u^{\rm NS}_{\varepsilon_{n}}(t_{n},u_{0})-u^{\rm E}(t_{n},u_{0})\right\|_{B^{s}_{p,\infty}}\geq c_{0}(2\pi)^{d-1}.$
@@
Acknowledgements
J. Li is supported by the National Natural Science Foundation of China (No.12161004), Innovative High end Talent Project in Ganpo Talent Program (No.gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No.20232BCJ23009).
@@ Acknowledgements
Data Availability No data was used for the research described in the article.
Conflict of interest The authors declare that they have no conflict of interest.

### v3

@@ 2 Preliminaries
$(\mathcal{F}u)(k)=\widehat{u}(k)=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}e^{-\mathrm{i}x\cdot k}u(x)\mathrm{d}x.$
@@
$u(x)=\sum_{k\in\mathbb{Z}^{d}}\widehat{u}(k)e^{\mathrm{i}x\cdot k}.$
We are interested in solutions which take values in the Besov space $B_{p,r}^{s}(\mathbb{T}^{d})$. Recall that one way to define this space requires a dyadic partition of unity. Given a smooth bump function $\chi$ supported on the ball of radius $4/3$, and equal to 1 on the ball of radius $3/4$, we set $\varphi(\xi)=\chi(2^{-1}\xi)-\chi(\xi)$ and $\varphi_{j}(\xi)=\varphi(2^{-j}\xi)$ and then we deduce that $\varphi$ satisfies that ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{T}^{d}:3/4\leq|\xi|\leq 8/3\right\}$ and $\varphi(\xi)\equiv 1$ for $4/3\leq|\xi|\leq 3/2$. Using this partition, we define the periodic dyadic blocks as follows
@@
$\Delta_{-1}u=\sum_{\xi\in\mathbb{Z}^{d}}\chi(\xi)\widehat{u}(\xi)e^{\mathrm{i}x\cdot\xi},$
$\Delta_{j}u=\sum_{\xi\in\mathbb{Z}^{d}}\varphi_{j}(\xi)\widehat{u}(\xi)e^{\mathrm{i}x\cdot\xi},\quad\text{ if }\quad j\geq 0.$
@@
Lemma 2.1.
Let $3\leq m\in\mathbb{Z}$ and $1\leq j\in\mathbb{Z}$, we have
$\Delta_{j}\left[\cos\left(\frac{11}{8}2^{m}x_{1}\right)\right]=\begin{cases}0,&\text{if}\quad j\neq m,\\ \cos\left(\frac{11}{8}2^{m}x_{1}\right),&\text{if}\quad j=m.\end{cases}$
Proof.
We set $\lambda=\frac{11}{8}2^{m}$ and $\vec{\lambda}=(\lambda,0,\cdots,0)$ for simplicity. Notice that $\varphi_{j}(k)$ is symmetric, i.e., $\varphi_{j}(k)=\varphi_{j}(|k|)$, we deduce
$\Delta_{j}\left(\cos(\lambda x_{1})\right)$ $=\frac{1}{2}\sum_{k\in\mathbb{Z}^{d}}\varphi_{j}(k)\left(\mathbf{1}_{\vec{\lambda}}(k)+\mathbf{1}_{-\vec{\lambda}}(k)\right)e^{\mathrm{i}x\cdot k}$
$=\frac{1}{2}\sum_{\ell\in\mathbb{Z}}\varphi_{j}(\ell)\left(\mathbf{1}_{\lambda}(\ell)+\mathbf{1}_{-\lambda}(\ell)\right)e^{\mathrm{i}x_{1}\ell}$
@@
Thus we obtain the desired result of Lemma 2.1. ∎
Remark 2.1.
For large fixed $m$, $\varphi_{j}\left(\frac{11}{8}2^{m}x_{1}\right)$ vanish except for $j=m$ due to the support condition of $\varphi$ and $\varphi(k)\equiv 1$ for $4/3\leq|k|\leq 3/2$.
@@
where $f(x)$ is a bounded real-valued periodic function of one variable with the following form
@@
$u(t,x)=\left(1,f\left(x_{1}-t\right),0,\ldots,0\right)$ (3.3)
@@
where $g(x)$ is a bounded real-valued period function of one variable with the following form
@@
Obviously, it is a period function which also satisfies the incompressible Euler equations (E) with initial condition $u_{0}(x)=(1,g(x_{1}),0,\cdots,0)$.
@@
$\Delta_{n}\left[f_{n}(t,x_{1})-f(x_{1}-t_{n})\right]=2^{-ns}\left[\cos\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{e}_{n}(t)+\sin\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{f}_{n}(t)\right]\mathbf{g}_{n}(t),$
@@
Taking the $\limsup$, we deduce that
$\limsup_{\varepsilon_{n}\to 0^{+}}\left\|u^{\rm NS}_{\varepsilon_{n}}(t_{n},u_{0})-u^{\rm E}(t_{n},u_{0})\right\|_{B^{s}_{p,\infty}}\geq c_{0}(2\pi)^{d-1}.$
@@
6 Proof of Theorems 1.5-1.6
Choice of initial data. Let us fix an $s>0$ and define the initial data which is a bounded real-valued periodic function of one variable with the following form
$\theta_{0}(\mathbf{x})=f(x_{1})=\sum\limits^{\infty}_{j=3}2^{-js}\cos\left(\frac{11}{8}2^{j}x_{1}\right),\quad\mathbf{x}=(x_{1},x_{2})\in\mathbb{T}^{2}.$
It is not difficult to check that $\theta_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ for any $1\leq p\leq\infty$.
Construction of solution. We are interested in solutions of the form
$\theta(t,\mathbf{x})=f\left(t,x_{1}\right),$
then it must have
$u\cdot\nabla\theta=u_{1}\partial_{1}\theta+u_{2}\partial_{2}\theta=0.$
In fact,
$\partial_{2}\theta=0\quad\text{and}\quad u_{1}=-\Lambda^{-1}\partial_{2}\theta=0.$
Thus $\theta(t,\mathbf{x})=e^{-\kappa\Lambda^{\alpha}t}\theta_{0}(x_{1})$ is a periodic solution of (QG). We shall compare the solution of (QG) with the initial data. Obviously,
$\theta(t,\mathbf{x})-\theta_{0}(\mathbf{x})$ $=\left(e^{-\kappa\Lambda^{\alpha}t}-1\right)\theta_{0}(\mathbf{x}).$
By Lemma 2.1, one has for some $n$ large enough
$\Delta_{j}\left(\left(e^{-\kappa\Lambda^{\alpha}t}-1\right)\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right)$ $=\begin{cases}0,&\text{if}\quad j\neq n,\\ \left(e^{-\kappa\left(\frac{11}{8}2^{n}\right)^{\alpha}t}-1\right)\cos\left(\frac{11}{8}2^{n}x_{1}\right),&\text{if}\quad j=n,\end{cases}$
which gives that
$\Delta_{n}\left[\theta(t,\mathbf{x})-\theta_{0}(\mathbf{x})\right]=2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right)\left[e^{-\kappa\left(\frac{11}{8}2^{n}\right)^{\alpha}t}-1\right].$ (6.1)
Proof of Theorem 1.5. Letting $\kappa\left(\frac{11}{8}2^{n}\right)^{\alpha}t_{n}=1$ in (6.1), one has
$\Delta_{n}\left[\theta(t_{n},\mathbf{x})-\theta_{0}(\mathbf{x})\right]$ $=(1-e^{-1})2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right).$
Combining the above, we have
$\|\theta(t_{n},\mathbf{x})-\theta_{0}(\mathbf{x})\|_{B^{s}_{p,\infty}(\mathbb{T}^{2})}$ $\geq 2^{ns}\left\|\Delta_{n}\left[\theta(t_{n},\mathbf{x})-\theta_{0}(\mathbf{x})\right]\right\|_{L^{p}(\mathbb{T}^{2})}$
$=\left(1-e^{-1}\right)\cdot\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}(\mathbb{T}^{2})}$
$=2\pi c_{0}\left(1-e^{-1}\right).$
Taking the $\limsup$, we deduce that
$\limsup_{t_{n}\to 0^{+}}\left\|\theta(t_{n},\mathbf{x})-\theta_{0}\right\|_{B^{s}_{p,\infty}}\geq\pi c_{0}.$
Thus we complete the proof of Theorem 1.5.
Proof of Theorem 1.6. We should notice that, the (QG) with $\kappa=0$ has a unique steady-state solution $\theta^{0}(t,\mathbf{x})=\theta_{0}(\mathbf{x})$. Let $\kappa_{n}=t_{n}=\left(\frac{11}{8}2^{n}\right)^{-\frac{\alpha}{2}}$ in (6.1), which gives that
$\Delta_{n}\left[\theta^{\kappa_{n}}(t_{n},\mathbf{x})-\theta^{0}(t_{n},\mathbf{x})\right]$ $=(1-e^{-1})2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right),$
and thus we deduce that
$\limsup_{\kappa_{n}\to 0^{+}}\left\|\theta^{\kappa_{n}}(t_{n},\mathbf{x})-\theta^{0}(t_{n},\mathbf{x})\right\|_{B^{s}_{p,\infty}}\geq\pi c_{0}.$
Notice that $t_{n}\to 0^{+}$ and $\kappa_{n}\to 0^{+}$ as $n\to\infty$, we complete the proof of Theorem 1.6.
7 Proof of Theorem 1.7
Lemma 7.1.
Let $s>0$. We define an initial data $\theta_{0}$ as follows
$\theta_{0}(\mathbf{x}):=\sum^{\infty}_{j=3}2^{-js}\sin\left(\frac{11}{8}2^{j}x_{1}\right)+\sin x_{2},\quad\mathbf{x}=(x_{1},x_{2})\in\mathbb{T}^{2}.$
Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ and some positive constants $C,c$ such that
$\left\|\theta_{0}\right\|_{B^{s}_{p,\infty}\cap L^{p}(\mathbb{T}^{2})}\leq C,$
$2^{ns}\left\|\Delta_{n}(u_{0}\cdot\nabla\theta_{0})\right\|_{L^{p}(\mathbb{T}^{2})}\geq c2^{n}.$
Proof.
The first is obvious. Notice that
$\Lambda^{-1}\cos(\lambda x_{i})=\frac{1}{2}\sum_{k\in\mathbb{Z}^{2}}\frac{1}{|k|}\left(\mathbf{1}_{\vec{\lambda}}(k)+\mathbf{1}_{-\vec{\lambda}}(k)\right)e^{\mathrm{i}x\cdot k}=\frac{1}{\lambda}\cos(\lambda x_{i}),\quad i=1,2,$
then
$u_{0}^{(1)}(x):=-\Lambda^{-1}\partial_{x_{2}}\theta_{0}(\mathbf{x})=-\cos x_{2},$
$u_{0}^{(2)}(x):=\Lambda^{-1}\partial_{x_{1}}\theta_{0}(\mathbf{x})=\sum^{\infty}_{j=3}2^{-js}\cos\left(\frac{11}{8}2^{j}x_{1}\right),$
and thus
$u_{0}\cdot\nabla\theta_{0}=\cos x_{2}\cdot\sum^{\infty}_{j=3}2^{-js}\left(1-\frac{11}{8}2^{j}\right)\cos\left(\frac{11}{8}2^{j}x_{1}\right).$
From which, we have
$\Delta_{n}(u_{0}\cdot\nabla\theta_{0})=\cos x_{2}\cdot 2^{-ns}\left(1-\frac{11}{8}2^{n}\right)\cos\left(\frac{11}{8}2^{n}x_{1}\right),$
which in turn gives
$2^{ns}\left\|\Delta_{n}(u_{0}\cdot\nabla\theta_{0})\right\|_{L^{p}(\mathbb{T}^{2})}$ $=\left(\frac{11}{8}2^{n}-1\right)\left\|\cos x_{2}\right\|_{L^{p}([0,2\pi])}\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}([0,2\pi])}$
$=\left(\frac{11}{8}2^{n}-1\right)\left(2\int^{\pi}_{0}|\cos x|^{p}\mathrm{d}x\right)^{2/{p}}\geq c2^{n},$
thus we complete the proof of Lemma 7.1. ∎
The following Lemma involves the nonlinear perturbation for the solution map of (QG) and is standard.
Lemma 7.2.
Assume that $s>0$ and $1\leq p\leq\infty$. Then we have for $t\in(0,T]$
$\left\|u\theta-u_{0}\theta_{0}\right\|_{\dot{B}^{s}_{p,\infty}}\leq C\left(\|\theta-\theta_{0}\|_{B^{s}_{p,\infty}}\|\theta_{0}\|_{L^{\infty}}+\|\theta-\theta_{0}\|_{L_{\mathcal{R}}^{\infty}}\|\theta_{0}\|_{B^{s}_{p,\infty}}\right).$ (7.1)
From (QG) and the Newton-Leibniz formula, it follows that
$\theta(t)-\theta_{0}$ $=-tu_{0}\cdot\nabla\theta_{0}-\int_{0}^{t}\mathord{{\rm div}}\left(u\theta-u_{0}\theta_{0}\right)\mathrm{d}\tau.$
Using the triangle inequality and Lemma 7.2, we have
$\left\|\theta(t)-\theta_{0}\right\|_{B^{s}_{p,\infty}}$ $\geq 2^{{ns}}\left\|\Delta_{n}\left(\theta(t)-\theta_{0}\right)\right\|_{L^{p}}$
$\geq t2^{{n}s}\left\|\Delta_{n}\big(u_{0}\cdot\nabla\theta_{0}\big)\right\|_{L^{p}}-t2^{n}2^{{n}s}\left\|\Delta_{n}(u\theta-u_{0}\theta_{0})\right\|_{L^{p}}$
$\geq t2^{n}\left(c-C\|\theta_{0}\|_{B^{s}_{p,\infty}\cap L^{\infty}}\|\theta(t)-\theta_{0}\|_{B^{s}_{p,\infty}\cap L_{\mathcal{R}}^{\infty}}\right).$
Suppose that (1.1) were not true for some $s>0$ and $1\leq p\leq\infty$. We pick $0<t_{0}\leq t$ such that for all $\tau\in[0,t_{0}]$
$\left\|\theta(t)-\theta_{0}\right\|_{B^{s}_{p,\infty}\cap L_{\mathcal{R}}^{\infty}}<\min\left(1,\frac{c}{2C\|\theta_{0}\|_{B^{s}_{p,\infty}\cap L^{\infty}}}\right).$
Thus, picking $t_{n}=M2^{-n}$ with large $M>0$, we have
$1>\left\|\theta(t_{n})-\theta_{0}\right\|_{B^{s}_{p,\infty}}\geq\frac{c}{2}M,$
which leads to a contraction for large $M$. This completes the proof of Theorem 1.7.
@@
Data Availability
No data was used for the research described in the article.
Conflict of interest
The authors declare that they have no conflict of interest.
Funding
J. Li is supported by the National Natural Science Foundation of China (12161004), Innovative High end Talent Project in Ganpo Talent Program (gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (20232BCJ23009), Jiangxi Provincial Natural Science Foundation (20252BAB22004). X.Wu is supported by the Natural Science Foundation of Henan Province (252300421774).